OpenStudy (anonymous):

Mustafa contributes 11% of his 67200 annual salary to his 401(k)plan what is his pretax income

OpenStudy (anonymous):

11*672=7392

OpenStudy (anonymous):

\[\frac{ 11 }{ 100}*67200=11*672=7392\]

OpenStudy (anonymous):

so the answer is

OpenStudy (anonymous):

obviously, thing that is written above lol 7392

OpenStudy (anonymous):

my choices are 14,993.85 , 71,41.92 ,49,016, 59,808

OpenStudy (anonymous):

B

OpenStudy (anonymous):

then m sure it is one of the four options mentioned above

OpenStudy (anonymous):

LOLLLLLLLLLL

OpenStudy (anonymous):

you know what i have been solving problems but i havent received one medal so far :(